OpenStudy (anonymous):

Write an equation of a line whose graph is parallel to the graph of y = 3x – 10.

OpenStudy (anonymous):

what is "line" in this case?

OpenStudy (anonymous):

A line on a graph I'm not sure

OpenStudy (campbell_st):

what is the slope of the given line..that is the important part of the question... your line is in the form y = mx + b m = slope and b = y-intercept so what is the value of m in the line you were given?

OpenStudy (anonymous):

I wasnt?

OpenStudy (campbell_st):

you have been given the line y = 3x -10 y = mx + b is the slope intercept form, where m = slope compare the 2 equations, what do you think the slope is..?

OpenStudy (anonymous):

ohh okay one sec

OpenStudy (anonymous):

@campbell_st I'm not sure can you please just walk me through it from the beginning

OpenStudy (campbell_st):

ok... in the line you are given, what number is in the m position..?

OpenStudy (anonymous):

3

OpenStudy (campbell_st):

great, thats the slope.... parallel lines have the same slope so your equation is y = 3x + b... now for your parallel line pick a number for b..anything... what would you choose..

OpenStudy (anonymous):

Any number?

OpenStudy (campbell_st):

well make it sensible... not 346551034

OpenStudy (anonymous):

okay 5 @campbell_st

OpenStudy (campbell_st):

great so that is b so your equation is y = 3x + 5 means the slope is 3 and where the line cuts the y axis is y = 5 the given line was y = 3x - 10 slope is 3 (so the 2 lines are parallel) and it cuts the y-axis y = -10 hope it makes sense